Mary has accepted a teaching job that pays ​$25,000 for the first year. According to the​ Teacher's Union, Mary will get guaranteed salary increases of 3% per year. If Mary plans to teach for 27 years, what will be her total salary​ earnings?
Answer by Boreal(15235)

Sum of a geometric series with ratio=1.03
S=A(1-r^n)/(1-r). n=27, r=1.03
S=25000(1-(1-0.03^27))/(1-(1.03)); the fraction is 40.709, but round at end
S=$1,017,740.84
Rough check. Without the increase in salary, $25,000 for 27 years is $675,000.